# spacing=nonfrench

### v1

@@ 3.3.3 Almost-Dirichlet Principle
Now let $\Gamma$ be a quasicircle and $p>1,\,s\in(0,1)$: we will say that $\mathcal{B}_{p,p}^{s}(\Omega)$ satisfies an almost-Dirichlet principle if there exists $C>0$ such that
@@
Maz’ya [29] has proven the almost-Dirichlet principle for the unit disk for all $p>1,s\in(0,1)$ and his proof has been simplified by [30]. Theorem 3.16 shows that if $\Gamma$ satisfies the hypothesis of this theorem, then the almost-Dirichlet principle holds.
3.3.4 The Case of pp–Dirichlet Energy
In this part we highlight the case of $p$-Dirichlet energy corresponding to $\omega\equiv 1$, i.e., s=11/ps=1-1/p.
Theorem 3.19.
If $\Gamma$ is a quasicircle then for all $p\in\left(2,\frac{2K_{*}-h(\Gamma)}{K_{*}-1}\right)$, there exists a constant $C=C(p,\Gamma,s)$ such that for $v\in C_{c}^{\infty}(\mathbb{C})$ ,
$\iint_{\Omega}|\nabla u(z)|^{p}dxdy\leq C\iint_{\Omega}|\nabla v(z)|^{p}dxdy,$ (3.10)
where $u$ stands for the harmonic extension of $v|_{\Gamma}$ in $\Omega$.
The proof is just by putting s=11/ps=1-1/p in Theorem 3.16 and taking care of the constraints: details are left to the reader. In particular, if $\Gamma$ is chord-arc then (3.10) holds for $p\in[2,3)$, and we have the following extension if $K<2$:
Corollary 3.20.
If $\Gamma$ is chord-arc with $K<2$ then the almost-Dirichlet principle holds for $2\leq p<\frac{2K-1}{K-1}.$
Notice that the corollary implies that for any $p>2$ the almost-Dirichlet principle is valid if $K$ is close enough to $1$.
4 Fractional Sobolev Spaces on Chord-arc Domains
4.1 A Review of Classical Facts about Fractional Sobolev Spaces on Plane Domains.
4.1.1 Fractional Sobolev Spaces on Plane Domains
We have already dealt with the Sobolev space $W^{1,p}(\mathbb{C})$, $1<p<\infty$ consisting of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{C})$ such that $f$ is locally integrable and $\nabla f$, taken in the sense of distributions, is in $L^{p}(\mathbb{C})$. The semi-norm
@@ 4.1.1 Fractional Sobolev Spaces on Plane Domains
equips $W^{1,p}(\mathbb{C})$ with a structure of Banach space modulo constants called homogeneous Sobolev space in the literature. The inhomogeneous Sobolev space is given by $\mathbb{W}^{1,p}(\mathbb{C})=W^{1,p}(\mathbb{C})\cap L^{p}(\mathbb{C})$. It is a Banach space equipped with the norm
@@
Let $\Omega$ be an open set in $\mathbb{C}$. $W^{1,p}(\Omega)$ and $\mathbb{W}^{1,p}(\Omega)$ should be defined similarly. We point out that if the open set $\Omega$ supports classical Poincaré inequality
@@
for every $f\in C_{c}^{\infty}(\Omega)$ where the constant $C$ depends only on $\Omega$ and $p$ then both spaces are known to coincide. For instance, the bounded chord-arc domain is a such domain. The bounded quasidisk also supports Poincaré inequality since it is an extension domain for Sobolev spaces $W^{1,p}(\mathbb{C})$ (see [22, 8]) so that one can apply Poincaré inequality in a bigger sufficiently regular domain which, combined with the extension theorem, leads to the conclusion.
For $0<s<1$ we define the fractional Besov-Sobolev space $B_{p,p}^{s}(\mathbb{C})$ as the space of $L^{p}(\mathbb{C})$-functions $f:\mathbb{C}\to\mathbb{C}$ such that
@@
We remark that suppose $s\geq 1$ and $\Omega$ is a connected open set in $\mathbb{C}$ then any measurable function $f:\Omega\to\mathbb{C}$ such that the integral of the above integrand $|f(z)-f(\zeta)|^{p}/|z-\zeta|^{2+sp}$ over $\Omega\times\Omega$ converges is actually constant. These spaces may be viewed as intermediate spaces between $W^{0,p}(\mathbb{C})$, i.e., $L^{p}(\mathbb{C})$ and $\mathbb{W}^{1,p}(\mathbb{C})$. More precisely, they may be obtained as real-interpolate spaces between these two spaces: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s,p}=B^{s}_{p,p}(\mathbb{C})$, see [40]. We call them Besov-Sobolev spaces because if one uses complex interpolation instead we obtain the spaces $H_{p}^{s}(\mathbb{C})$ consisting of Bessel potential of $L^{p}$-functions: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s}=H^{s}_{p}(\mathbb{C})$. and $B_{p,p}^{s}(\mathbb{C})=H_{p}^{s}(\mathbb{C})$ only if $p=2$ (if $p\neq 2$ these two spaces are not even isomorphic).
For further use we extend the definition of Besov-Sobolev spaces to $s\in(1,2)$ by saying that $f\in B_{p,p}^{s}(\mathbb{C})$ with $s\in(1,2)$ if $f\in B_{p,p}^{s-1}(\mathbb{C})$ and $\nabla f\in B_{p,p}^{s-1}(\mathbb{C})$. Similarly, it can be further extended to other non-integer orders $s>0$. It is known that the space $C_{c}^{\infty}(\mathbb{C})$ is dense in $B_{p,p}^{s}(\mathbb{C})$ for non-integer orders $s>0$ and $1<p<\infty$.
@@
We define $B_{p,p}^{s}(\Omega)$ simply as the space of restrictions to $\Omega$ of functions in $B_{p,p}^{s}(\mathbb{C})$. Such a function is in particular such that
@@
4.1.2 Fractional Sobolev Spaces on Chord-arc Curves
If $\Omega$ is a (bounded or unbounded) chord-arc domain with boundary $\Gamma$ we may define for $0<s<1$, $p>1$,
@@ 4.1.2 Fractional Sobolev Spaces on Chord-arc Curves
We define the trace operator $\gamma$ on $C_{c}^{\infty}(\mathbb{C})$ by $\gamma(u)=u|_{\Gamma}$.
@@
Let $s\in(1/p,1+1/p)$. The trace operator $\gamma$ can be extended to a bounded linear operator from $B_{p,p}^{s}(\Omega)$ onto $B_{p,p}^{s-1/p}(\Gamma)$, the order having $1/p$-loss, whose kernel is the closure of $C_{c}^{\infty}(\Omega)$ in $B_{p,p}^{s}(\Omega)$.
4.2 Plemelj-Calderón Problem on Chord-arc Domains
We now come back to Plemelj-Calderón problem. By a deep theorem of David [11] using the Lipschitz result [10], the Cauchy integral operator $T$ is bounded on $L^{p}(\Gamma)$ (i.e., $B_{p,p}^{0}(\Gamma)$) for every $p\in(1+\infty)$ (see section 1.3). Using it we will show the boundedness of the operator $T$ on $B_{p,p}^{1}(\Gamma)$.
@@ 4.2 Plemelj-Calderón Problem on Chord-arc Domains
The operator $T$ is bounded on the Sobolev space $B_{p,p}^{1}(\Gamma)$, and moreover, the operator norm on $B_{p,p}^{1}(\Gamma)$ equals to that on $L^{p}(\Gamma)$.
Before we proceed to the proof of this theorem, one must properly define what we mean by $B^{1}_{p,p}(\Gamma)$ when $\Gamma$ is a chord-arc curve (or $K$-chord-arc curve for specifying its chord-arc constant). Without loss of generality we suppose $\text{length}(\Gamma)=2\pi$. Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map from the unit disk $\mathbb{D}$ onto the interior domain $\Omega$ of $\Gamma$, which can be extended to a homeomorphism of the closures such that $\varphi$ restricted on $\mathbb{T}$ is a quasisymmetry (see [35]). Let $t\mapsto z(t)$ denote the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. It is easy to see $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. Indeed, for any $t_{1},t_{2}\in[0,2\pi)$,
@@
To prepare for the proof we introduce the following definition. Set $\Gamma_{r}=\tau(|z|=r)$ to be "circular curves", where $\tau$ is a Riemann map that takes $\mathbb{D}$ onto $\Omega$. The Hardy space $E^{p}(\Omega)$ of index $p$ on the chord-arc domain $\Omega$ consists of holomorphic functions $F$ on $\Omega$ with a finite norm
@@
on $\Gamma$. Here, $G_{i}$ and $G_{e}$ belong to the Hardy spaces of index $p$ on $\Omega_{i}$ and $\Omega_{e}$, respectively, and then $G_{i}$ and $G_{e}$ have non-tangential boundary limits almost everywhere on $\Gamma$ with respect to the arc-length measure, denoted still by $G_{i}$ and $G_{e}$, so that $\|G_{i}\|_{L^{p}(\Gamma)}$ and $\|G_{e}\|_{L^{p}(\Gamma)}$ are both controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
Recall that $G_{i}\in E^{p}(\Omega_{i})\subset E^{1}(\Omega_{i})$, so that $G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. Let now $\Phi_{i}$ be an anti-derivative of $G_{i}$ on $\Omega_{i}$. Then we see $(\Phi_{i}\circ\tau)^{\prime}=G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. By a result of Hardy-Littlewood (see e.g. p.89 in [16]), we conclude that $\Phi_{i}\circ\tau$ is continuous on $\overline{\mathbb{D}}_{i}$ and absolutely continuous on $\mathbb{T}$ such that $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=\lim_{r\to 1}(\Phi_{i}\circ\tau)^{\prime}(r\zeta)$ almost everywhere on $\mathbb{T}$, and thus $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=G_{i}\circ\tau(\zeta)\tau^{\prime}(\zeta)$ almost everywhere on $\mathbb{T}$. Using it, let us define $\Phi_{i}^{\prime}=G_{i}$ on $\Gamma$, and we can similarly define $\Phi_{e}^{\prime}=G_{e}$ on $\Gamma$. Combined with (4.1) that leads to, by adjusting the constants,
@@
on $\Gamma$ with the norm of $\Phi_{i,e}$ in $B^{1}_{p,p}(\Gamma)$ controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$. By Plemelj formula we conclude that the operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$ with respect to the norm $\|\cdot\|_{B_{p,p}^{1}(\Gamma)}$, and moreover, we have $\|T\|_{B_{p,p}^{1}(\Gamma)\to B_{p,p}^{1}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$. ∎
Now we get ready to show the main result of this section: the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for $0<s<1$. Because of the isomorphism between $B_{p,p}^{s}(\Gamma)$ and $B_{p,p}^{s}(\mathbb{T})$, this boils down to proving that the operator with kernel
@@
The following functorial property of complex interpolation is a basic assertion in interpolation theory. Let $\{A_{0},A_{1}\}$ and $\{B_{0},B_{1}\}$ be two interpolation couples of Banach spaces and let $L$ be a linear operator mapping from $A_{0}+A_{1}$ into $B_{0}+B_{1}$ such that its restriction to $A_{j}$ is a linear and bounded operator from $A_{j}$ into $B_{j}$ with norm $M_{j}$, where $j=0,1$. Then the restriction of $L$ to $A_{0},A_{1}]_{\theta}$, $0<\theta<1$ is a linear and bounded operator from $A_{0},A_{1}]_{\theta}$ into $B_{0},B_{1}]_{\theta}$ with norm $M_{\theta}$. Further, it is known that the complex interpolation method is an exact interpolation functor of exponent $\theta$ (see e.g. p.88 in [4]) meaning that
@@
By taking $s_{0}=0$ and $s_{1}=1$ in Theorem 4.3, we have that
@@
It follows from Plemelj formula that every function $f\in B_{p,p}^{s}(\Gamma)$ may be written uniquely as $f=\Phi_{i}+\Phi_{e}$ with $\|\Phi_{i,e}\|_{B_{p,p}^{s}(\Gamma)}\leq C\|f\|_{B_{p,p}^{s}(\Gamma)}$ for some constant $C$, $\Phi_{i,e}$ being boundary values of holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively. The uniqueness of decomposition for $0<s\leq 1$ just follows from the case of $s=0$.
4.2.1 On a Theorem of Murai
Clearly, Theorem 4.4 is also valid for chord-arc curves $\Gamma$ passing through $\infty$, and in particular for Lipschitz curves $\Gamma$ (i.e., there is a Lipschitz function $A:\mathbb{R}\to\mathbb{R}$ whose graph is $\Gamma$). A theorem by Murai [31] states that for a Lipschitz curve $\Gamma$ with Lipschitz norm $M$, $\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$ is no more than $C(1+M)^{3/2}$ (see also [12]), where $C$ is a universal constant. If we plug this information in the preceding theorem we obtain the same bound for $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$. But the norm $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$ depends only on the $L^{p}(\mathbb{C})$-boundedness of the Beurling transform (see section 2), so that it is actually independent of $M$. In order to get better estimates of $\|T\|_{B_{p,p}^{s}(\Gamma)\to B_{p,p}^{s}(\Gamma)}$ we thus use Calderón’s interpolation theorem between $L^{p}(\Gamma)$ and $B_{p,p}^{1-1/p}(\Gamma)$ for 0<s<11/p0<s<1-1/p and between $B_{p,p}^{1-1/p}(\Gamma)$ and $B_{p,p}^{1}(\Gamma)$ for 11/p<s<11-1/p<s<1. We use more precisely that

### v2

@@ 3.4 Almost-Dirichlet principle
Now, let $\Gamma$ be a quasicircle, and $p>1,\,s\in(0,1)$. We will say that $\mathcal{B}_{p,p}^{s}(\Omega)$ satisfies an almost-Dirichlet principle if there exists $C>0$ such that
@@
Maz’ya [32] has proven the almost-Dirichlet principle for the unit disk for all $p>1,s\in(0,1)$ and his proof has been simplified by [33]. Theorem 3.5 shows that if $\Gamma$ satisfies the hypotheses of this theorem, then the almost-Dirichlet principle holds.
4 Plemelj-Calderón problem for $B_{p,p}^{s}(\Gamma)$ on chord-arc curves: the rectifiable case
4.1 A review of classical facts about fractional Sobolev spaces in plane domains.
We have already dealt with the Sobolev space $W^{1,p}(\mathbb{C})$, $1<p<\infty$, consisting of tempered distributions $f$ such that $f$ is locally integrable and $\nabla f$, taken in the sense of distributions, is in $L^{p}(\mathbb{C})$. The semi-norm
@@ 4 Plemelj-Calderón problem for $B_{p,p}^{s}(\Gamma)$ on chord-arc curves: the rectifiable case
equips $W^{1,p}(\mathbb{C})$ with a structure of Banach space modulo constants called the homogeneous Sobolev space in the literature. The inhomogeneous Sobolev space is given by $\mathbb{W}^{1,p}(\mathbb{C})=W^{1,p}(\mathbb{C})\cap L^{p}(\mathbb{C})$. It is a Banach space equipped with the norm
@@
Let $\Omega$ be an open set in $\mathbb{C}$. $W^{1,p}(\Omega)$ and $\mathbb{W}^{1,p}(\Omega)$ should be defined similarly. We point out that if the open set $\Omega$ satisfies the classical Poincaré inequality, namely,
@@
for every $f\in C_{c}^{\infty}(\Omega)$ where the constant $C$ depends only on $\Omega$ and $p$ then both spaces coincide. For instance, bounded chord-arc domains are such domains. The bounded quasidisks also satisfy Poincaré inequality, since it is an extension domain for Sobolev spaces $W^{1,p}(\mathbb{C})$ (see [22, 8]) so that one can apply Poincaré inequality in a sufficiently regular larger domain, which, combined with the extension theorem, leads to the conclusion.
For $0<s<1$ we define the fractional Besov-Sobolev space $B_{p,p}^{s}(\mathbb{C})$ in $\mathbb{C}$ as the space of $L^{p}(\mathbb{C})$-functions $f:\mathbb{C}\to\mathbb{C}$ such that
@@
(We remark that if $s\geq 1$ and $\Omega$ is a connected open set in $\mathbb{C}$ then any measurable function $f:\Omega\to\mathbb{C}$ such that the integral of the above integrand $|f(z)-f(\zeta)|^{p}/|z-\zeta|^{2+sp}$ over $\Omega\times\Omega$ converges is actually constant.) These spaces may be viewed as intermediate spaces between $W^{0,p}(\mathbb{C})$, i.e., $L^{p}(\mathbb{C})$ and $\mathbb{W}^{1,p}(\mathbb{C})$. More precisely, they may be obtained as real-interpolate spaces between these two spaces: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s,p}=B^{s}_{p,p}(\mathbb{C})$, see [46]. Notice that they are also special Besov spaces. We call them fractional Besov-Sobolev spaces. If we use complex interpolation instead, we obtain the spaces $H_{p}^{s}(\mathbb{C})$ consisting of Bessel potentials of $L^{p}$-functions: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s}=H^{s}_{p}(\mathbb{C})$. It is known that $B_{p,p}^{s}(\mathbb{C})=H_{p}^{s}(\mathbb{C})$ only if $p=2$ (if $p\neq 2$ these two spaces are not even isomorphic).
For further use, we extend the definition of fractional Besov-Sobolev spaces to the case of $s\in(1,2)$ by saying that $f\in B_{p,p}^{s}(\mathbb{C})$ with $s\in(1,2)$ if $f\in B_{p,p}^{s-1}(\mathbb{C})$ and $\nabla f\in B_{p,p}^{s-1}(\mathbb{C})$. Similarly, it can be extended to the case of other non-integer orders $s>0$. It is known that the space $C_{c}^{\infty}(\mathbb{C})$ is dense in $B_{p,p}^{s}(\mathbb{C})$ for non-integer orders $s>0$ and $1<p<\infty$.
@@
We define $B_{p,p}^{s}(\Omega)$ simply as the space of restrictions to $\Omega$ of functions in $B_{p,p}^{s}(\mathbb{C})$. Such a function , in particular, satisfies
@@
If $\Omega$ is a chord-arc domain with boundary $\Gamma$ we may define the fractional Besov-Sobolev spaces $B_{p,p}^{s}(\Gamma)$ on $\Gamma$ as follows: $0<s<1$, $p>1$,
@@
We define the trace operator $\gamma$ in $C_{c}^{\infty}(\mathbb{C})$ by $\gamma(u)=u|_{\Gamma}$.
@@
Let $\Omega$ be a chord-arc domain and let $s\in(1/p,1+1/p)$. The trace operator $\gamma$ can be extended to a bounded linear operator from $B_{p,p}^{s}(\Omega)$ onto $B_{p,p}^{s-1/p}(\Gamma)$, the order having $1/p$-loss, whose kernel is the closure of $C_{c}^{\infty}(\Omega)$ in $B_{p,p}^{s}(\Omega)$.
4.2 Plemelj-Calderón problem on chord-arc curves
We now come back to Plemelj-Calderón problem. Recall that the Cauchy integral operator $T$ is bounded on $L^{p}(\Gamma)$ (i.e., $B_{p,p}^{0}(\Gamma)$) for every $p\in(1+\infty)$ (see subsection 1.3) if $\Gamma$ is chord-arc. Our goal in this subsection is to show that the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for all $p\in(1,+\infty)$ and $s\in(0,1]$. We begin by proofing that the result is valid for $B_{p,p}^{1}(\Gamma)$ using that for $B_{p,p}^{0}(\Gamma)$.
@@ 4.2 Plemelj-Calderón problem on chord-arc curves
The operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$, and moreover, the operator norm on $B_{p,p}^{1}(\Gamma)$ is equal to that on $L^{p}(\Gamma)$.
Before we proceed to the proof of this theorem, one must properly define what we mean by $B^{1}_{p,p}(\Gamma)$ when $\Gamma$ is a chord-arc curve (or $K$-chord-arc curve to specify its chord-arc constant). Without loss of generality, we suppose $\text{length}(\Gamma)=2\pi$. Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map from the unit disk $\mathbb{D}$ onto the interior domain $\Omega$ of $\Gamma$, which can be extended to a homeomorphism of the closures such that $\varphi$ restricted on $\mathbb{T}$ is a quasisymmetry (see [39]). Let $t\mapsto z(t)$ denote the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. It is easy to see that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. Indeed, for any $t_{1},t_{2}\in[0,2\pi)$,
@@
To prepare for the proof, we introduce the following definition. Set $\Gamma_{r}=\tau(|z|=r)$ as "circular curves", where $\tau$ is a Riemann map that takes $\mathbb{D}$ to $\Omega$. The Hardy space $E^{p}(\Omega)$ of the index $p$ in the chord-arc domain $\Omega$ consists of holomorphic functions $F$ in $\Omega$ with finite norm
@@
on $\Gamma$. Here, $G_{i}$ and $G_{e}$ belong to the Hardy spaces of index $p$ in $\Omega_{i}$ and $\Omega_{e}$, respectively, and then $G_{i}$ and $G_{e}$ have non-tangential boundary limits almost everywhere on $\Gamma$ with respect to the arc-length measure, denoted still by $G_{i}$ and $G_{e}$, so that $\|G_{i}\|_{L^{p}(\Gamma)}$ and $\|G_{e}\|_{L^{p}(\Gamma)}$ are both controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
Recall that $G_{i}\in E^{p}(\Omega_{i})\subset E^{1}(\Omega_{i})$, so that $G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. Let now $\Phi_{i}$ be an anti-derivative of $G_{i}$ in $\Omega_{i}$. Then we see $(\Phi_{i}\circ\tau)^{\prime}=G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. By a result of Hardy-Littlewood (see e.g. p.89 in [16]), we conclude that $\Phi_{i}\circ\tau$ is continuous on $\overline{\mathbb{D}}_{i}$ and absolutely continuous on $\mathbb{T}$ such that $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=\lim_{r\to 1}(\Phi_{i}\circ\tau)^{\prime}(r\zeta)$ almost everywhere on $\mathbb{T}$, and thus $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=G_{i}\circ\tau(\zeta)\tau^{\prime}(\zeta)$ almost everywhere on $\mathbb{T}$. Using it, let us define $\Phi_{i}^{\prime}=G_{i}$ on $\Gamma$, and we can similarly define $\Phi_{e}^{\prime}=G_{e}$ on $\Gamma$. Combined with (4.1) that leads to, by adjusting the constants,
@@
on $\Gamma$ with the norm of $\Phi_{i,e}$ in $B^{1}_{p,p}(\Gamma)$ controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$. By Plemelj formulae we conclude that the operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$ with respect to the norm $\|\cdot\|_{B_{p,p}^{1}(\Gamma)}$, and moreover, we have $\|T\|_{B_{p,p}^{1}(\Gamma)\to B_{p,p}^{1}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$. ∎
Now we get ready to show the main result of this subsection: the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for $0<s<1$. Due to the isomorphism between $B_{p,p}^{s}(\Gamma)$ and $B_{p,p}^{s}(\mathbb{T})$, this boils down to proving that the operator with kernel
@@
The following functorial property of complex interpolation is a basic assertion in interpolation theory. Let $(A_{0},A_{1})$ and $(B_{0},B_{1})$ be two interpolation couples of Banach spaces, and let $L$ be a linear operator mapping from $A_{0}+A_{1}$ to $B_{0}+B_{1}$ such that its restriction to $A_{j}$ is a linear and bounded operator from $A_{j}$ to $B_{j}$ with the norm $M_{j}$, where $j=0,1$. Then the restriction of $L$ to $A_{0},A_{1}]_{\theta}$, $0<\theta<1$, is a linear and bounded operator from $A_{0},A_{1}]_{\theta}$ to $B_{0},B_{1}]_{\theta}$ with the norm $M_{\theta}$. Furthermore, it is known that the complex interpolation method is an exact interpolation functor of exponent $\theta$ (see, e.g., p.88 in [4]) meaning that
@@
Taking $s_{0}=0$ and $s_{1}=1$ in Theorem 4.3, we have
@@
It follows from Plemelj formulae that every function $f\in B_{p,p}^{s}(\Gamma)$ can be written uniquely as $f=\Phi_{i}+\Phi_{e}$ with $\|\Phi_{i,e}\|_{B_{p,p}^{s}(\Gamma)}\leq C\|f\|_{B_{p,p}^{s}(\Gamma)}$ for some constant $C$, $\Phi_{i,e}$ being the boundary values of the holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively. The uniqueness of the decomposition for $0<s\leq 1$ follows from the case of $s=0$.
4.3 Plemelj-Calderón problem on Lipschitz curves
Clearly, Theorem 4.4 is also valid for chord-arc curves $\Gamma$ passing through $\infty$, and in particular for Lipschitz curves $\Gamma$ (i.e., there is a Lipschitz function $A:\mathbb{R}\to\mathbb{R}$ whose graph is $\Gamma$). A theorem by Murai [35] states that for a Lipschitz curve $\Gamma$ with the Lipschitz norm $M$, $\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$ is no more than $C(1+M)^{3/2}$ (see also [12]), where $C$ is a universal constant. If we plug this information into the preceding theorem, we obtain the same bound for $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$. But the norm $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$ depends only on the $L^{p}(\mathbb{C})$-boundedness of the Beurling transform (see Section 2), so that it is actually independent of $M$. In order to get better estimates of $\|T\|_{B_{p,p}^{s}(\Gamma)\to B_{p,p}^{s}(\Gamma)}$ we thus use Calderón’s interpolation theorem between $L^{p}(\Gamma)$ and $B_{p,p}^{1-1/p}(\Gamma)$ for 0<s<11/p0<s<1-1/p and between $B_{p,p}^{1-1/p}(\Gamma)$ and $B_{p,p}^{1}(\Gamma)$ for 11/p<s<11-1/p<s<1. We use more precisely that